OpenStudy (18jonea):

8 – 4x^3 + 4x^6 = 0, state the number of complex roots, the possible number of real roots,and the possible rational roots. @mathmale

OpenStudy (18jonea):

I know there are 6 roots possible and 0 ,2 ,4 , or 6 real roots possible now i just need to find the possible rational roots

OpenStudy (18jonea):

@Michele_Laino

OpenStudy (18jonea):

@TheSmartOne

OpenStudy (18jonea):

@phi

OpenStudy (phi):

If you put the expression in standard order: \[ 4x^6 – 4x^3+8= 0 \] and count the number of sign changes you get 2. that means there are either 2 or 0 positive real roots if you put in -1 for x you get 4 +4 + 8 =0 and there are no sign reversals, so no negative real roots.

OpenStudy (18jonea):

the possible answers are 1/8 1/4 1/2 1 2 4 or 1/4 1/2 1 2 4 8

OpenStudy (phi):

the highest exponent is 6, so 6 complex roots and either 2 or 0 real roots

OpenStudy (18jonea):

ok i got that part but i have to find the possible rational roots

OpenStudy (phi):

factors of 8 divided by factors of 4

OpenStudy (phi):

the biggest possible rational root will be 8/1 = 8 the smallest 1/4

OpenStudy (18jonea):

factors of 8 8 ,4,2,1 factors of 4 4,2,1

OpenStudy (18jonea):

ok so 1/4 1/2 1 2 4 and 8?

OpenStudy (phi):

yes

OpenStudy (18jonea):

Thank you so much!!!!!!!